The Geology Club of Gilroy Middle School took a field trip to Nisene Marks State Park to visit the epicenter of the 1989 earthquake. There were 3 adults for every 8 students on the field trip. A total of 44 people went on the field trip. How many adults were on the field trip? (Hint: a table may help)
Found 2 solutions by lwsshak3, lynnlo:
Answer by lwsshak3(11628)

The Geology Club of Gilroy Middle School took a field trip to Nisene Marks State Park to visit the epicenter of the 1989 earthquake. There were 3 adults for every 8 students on the field trip. A total of 44 people went on the field trip. How many adults were on the field trip?
**
let x=number of adults who went on the field trip
44-x=number of students who went on the field trip
..
3/8=x/(44-x)
cross-multiply
8x=3(44-x)
8x=132-3x
11x=132
x=132/11=12
number of adults who went on the field trip=12

Answer by lynnlo(4176)